# Jacobi polynomials for first-order generalized Reed–Muller codes

### 22 Dec 2023 v3

@@ 1. Introduction
Let $C=RM_{q}(1,m)$ and $C_{\ell}:=\{c\in C\mid\mathrm{wt}(c)=\ell\}$. In this paper, we call $C_{\ell}$ a shell of the code $C$ whenever it is non-empty. We show shells of $C$ support combinatorial $2$-designs but do not support combinatorial $3$-designs by using Jacobi polynomials. More precisely, the set $\mathcal{B}(C_{\ell}):=\{\mathrm{supp}(x)\mid x\in C_{\ell}\}$ forms the set of a combinatorial $2$-design but does not form a $3$-design. Herein, we always assume that a combinatorial $t$-design allows the existence of repeated blocks, and we exclude the trivial design $\mathcal{D}=(\Omega,\mathcal{B})$ where $\Omega=\{1,\dots,n\}$ and ={Ω,,$\mathcal{B}=\{\Omega,\dots,\Omega\}$.
@@
In [7], Miezaki and Nakasora provided the first non-trivial examples of a code that supports combinatorial $t$-designs for all weights and that supports $t^{\prime}$-designs for some weights with some $t^{\prime}>t$. Therefore, it is important to determine the value $t$ such that all shells of a code support a $t$-design, and no shell supports a $t+1$-design.
@@
Using this theorem, we show that the shells of $RM_{q}(1,m)$ and $RM_{q}(1,m)^{\perp}$ support $2$-designs.
@@
Let $C=RM_{q}(1,m)$. If $q\geq 3$, then for any $\ell\in\mathbb{N}$, $C_{\ell}$ is not a combinatorial $3$-design. Similarly, $(C^{\perp})_{\ell}$ is not a combinatorial $3$-design.
Using Theorem 1.3, we show that $C_{\ell}$ supports a $3$-$(v,k,(\lambda_{1},\lambda_{2}))$-design (see Section 2.1).
@@
By this theorem, we show that $C_{\ell}$ supports a combinatorial $4$-$(v,k,(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}))$-design.
@@
This paper is organized as follows. In Section 2, we define and give some basic properties of the codes, combinatorial $t$-designs, and Jacobi polynomials used in this paper. In Sections 3, 4, and 5, we show Theorems 1.1, 1.3, and 1.6, respectively.
@@ 2. Preliminaries
Let $q$ be a prime power. A $q$-ary linear code $C$ of length $n$ is a linear subspace of $\mathbb{F}_{q}^{n}$. The dual code of $C$ is the set of vectors which are orthogonal to any codewords in $C$: $C^{\perp}:=\{x\in\mathbb{F}_{q}^{n}\mid x\cdot c=0\ \mathrm{for\ all\ }c\in C\}$. For $c\in\mathbb{F}_{q}^{n}$, the weight $\mathrm{wt}(c)$ is the number of its nonzero components. The shell of a $-\ell$ codeword is the set of codewords whose weight is $\ell$: $C_{\ell}:=\{c\in C\mid\mathrm{wt}(c)=\ell\}$.
A combinatorial $t$-design is a pair $\mathcal{D}=(\Omega,\mathcal{B})$, where $\Omega$ is a set of points of cardinality $v$, and $\mathcal{B}$ is a collection of $k$-element subsets of $\Omega$ called blocks, with the property that any $t$-element subset of $\Omega$ is contained in precisely $\lambda$ blocks. Recall [4] for the definitions of various types of designs. A pair $\mathcal{D}=(\Omega,\mathcal{B})$ is a combinatorial design with parameters $t$-$(v,k,(\lambda_{1},\dots,\lambda_{N}))$ if $\mathcal{B}$ is a collection of $k$-element subsets of $\Omega$ called blocks and every $t$-element subset of $\Omega$ is contained in $\lambda_{i}$ blocks. Note that for $N=1$, the design coincides with an exactly $t$-design.
The support of a vector $x:=(x_{1},\dots,x_{n})$, $x_{i}\in\mathbb{F}_{q}$, is the set of indices of its nonzero coordinates: $\mathrm{supp}(x)=\{(c)\mid x_{i}\neq 0\}$. Let $\Omega:=\{1,\dots,n\}$ and $\mathcal{B}(C_{\ell}):=\{\mathrm{supp}(x)\mid x\in C_{\ell}\}$. Then for a code $C$ of length $n$, we say that the shell $C_{\ell}$ is a combinatorial $t$-design if $(\Omega,\mathcal{B}(C_{\ell}))$ is a combinatorial $t$-design. Similarly, we say that the shell $C_{\ell}$ is a combinatorial $t$-$(v,k,(\lambda_{1},\dots,\lambda_{N}))$-design if $(\Omega,\mathcal{B}(C_{\ell}))$ is a combinatorial $t$-$(v,k,(\lambda_{1},\dots,\lambda_{N}))$-design.
@@
This fact is useful because it suffices to consider the case that an element of $T$ is zero.
@@ 2.3. Notation
We next introduce some notation. Let $C=RM_{q}(1,m)$, $T\subset V$, and $t=|T|$. For $c\in C$, $i\in\{0,1,\dots,t\}$, $j\in\mathbb{F}_{q}$, we define
@@
$a_{i}:=|\{c\in C\setminus\mathbb{F}_{q}\mathbf{1}\mid\mathrm{wt}(c|_{T})=i\}|=a_{t-i,0},$
@@
$a_{i,0}$ =|$=|\{c\in C\setminus\mathbb{F}_{q}\mathbf{1}\mid n_{0,T}(c)=i\}|$
@@
Thus, we give coefficients of the Jacobi polynomial by (2.1). ∎
@@
Because $u_{1}$ and $u_{2}$ are not equal to 00 or each other, there exist $a,b\in\mathbb{F}_{q}\setminus\{0\}$ such that $a\neq b$ and {[$\{{}^{t}[a\ b]\}$ is a basis of $\mathrm{Im}\,A$. Hence, $\mathrm{rank}\,A=1$ and for any $v\in\mathrm{Im}\,A$,
@@
Then, we show that the shells of $RM_{q}(1,m)$ and $RM_{q}(1,m)^{\perp}$ do not support $3$-designs if $q\geq 3$.
@@
Let $C=RM_{q}(1,m)$. We give a proof relying on the properties of Jacobi polynomials. Let $T_{1}=\{0,u_{1},u_{2}\}\in\binom{V}{3}$, $T_{2}=\{0,v_{1},v_{2}\}\in\binom{V}{3}$, $A_{1}={}^{t}[u_{1}\ u_{2}]$, and $A_{2}={}^{t}[v_{1}\ v_{2}]$. We assume that $\mathrm{rank}\,A_{1}=2$ and $\mathrm{rank}\,A_{2}=1$. Indeed, if $q\geq 3$, there exist such $T_{1},T_{2}$. By Theorem 1.3,
@@
Since the coefficient of $z^{3}x^{q^{m-1}-\ell}y^{\ell-3}$ in $J_{C,T_{1}},J_{C,T_{2}}$ is non-zero whenever $C_{\ell}$ is non-empty, $C_{\ell}$ is not a $3$-design. By using Theorem 2.1, we have
@@
Thus, $(C^{\perp})_{\ell}$ is not a $3$-design. ∎
Finally, we claim that the shells of $RM_{q}(1,m)$ support $3$-$(v,k,(\lambda_{1},\lambda_{2}))$-designs.
@@
$b_{i,j}$ =|$=|\{c\in V\mid i\text{ elements of $Ac$ are equal to }j\}|$
$=q^{m-\mathrm{rank}\,A}\times|\{v\in\mathrm{Im}\,A\mid i\text{ elements of $v$ are equal to }j\}|,$
@@
$b_{i,0}$ =|$=|\{c\in V\mid i-1\text{ elements of $Ac$ are equal to }0\}|$
$=q^{m-\mathrm{rank}\,A}\times|\{v\in\mathrm{Im}\,A\mid i-1\text{ elements of $v$ are equal to }0\}|.$
@@
If $\mathrm{rank}\,A=3$, $A$ is a surjective linear map from $V$ to $\mathbb{F}_{q}^{3}$. Then, for all $v\in\mathbb{F}_{q}^{3}$, $|A^{-1}(v)|=q^{m-3}$. Thus,
@@
Before claiming the lemma under the condition of $\mathrm{rank}\,A=2$, we give a basis of $\mathrm{Im}\,A$. Since $\mathrm{Im}\,A$ is invariant under the right multiplication of an invertible matrix to $A$, we confine a basis of $\mathrm{Im}\,A$ to the following:
@@
If $\mathrm{rank}\,A=2$, we have the following equations.
@@
|$|\{(c_{1},c_{2})\in\mathbb{F}_{q}^{2}\mid c_{1}=j,c_{1}a+c_{2}b=j\}|$ =$=\left\{\begin{array}[]{ll}1&\text{if }b\neq 0\\ 0&\text{if }b=0\end{array}\right.,$
|$|\{(c_{1},c_{2})\in\mathbb{F}_{q}^{2}\mid c_{2}=j,c_{1}a+c_{2}b=j\}|$ =$=\left\{\begin{array}[]{ll}1&\text{if }a\neq 0\\ 0&\text{if }a=0\end{array}\right.,$
|$|\{(c_{1},c_{2})\in\mathbb{F}_{q}^{2}\mid c_{1}=j,c_{2}=j,c_{1}a+c_{2}b=j\}|$ =$=\left\{\begin{array}[]{ll}1&\text{if }a+b=1\\ 0&\text{otherwise}\end{array}\right..$
@@
Second, we give $b_{1}$. Similarly,
$b_{1,0}=\left\{\begin{array}[]{ll}(q-1)^{2}&\text{if }ab=0,\\ (q-1)(q-2)&\text{otherwise}.\end{array}\right.$
@@
Finally, by using Lemma 2.4, we give $a_{0},\dots,a_{4}$ as follows:

### 25 Feb 2024 v4

# Jacobi polynomials for the first-order generalized Reed–Muller codes

@@ 1. Introduction
Let $C=RM_{q}(1,m)$ and $C_{\ell}:=\{c\in C\mid\mathrm{wt}(c)=\ell\}$. In this paper, we call $C_{\ell}$ a shell of the code $C$ whenever it is non-empty. We show shells of $C$ are combinatorial $2$-designs but are not combinatorial $3$-designs by using Jacobi polynomials. More precisely, the set $\mathcal{B}(C_{\ell}):=\{\mathrm{supp}(x)\mid x\in C_{\ell}\}$ forms the set of blocks of a combinatorial $2$-design but does not form a $3$-design. Herein, we always assume that a combinatorial $t$-design allows the existence of repeated blocks, and we exclude the trivial design $\mathcal{D}=(\Omega,\mathcal{B})$ where $\Omega=\{1,\dots,n\}$ and ={Ω,,$\mathcal{B}=\{\Omega,\dots,\Omega\}$.
@@
In [7], Miezaki and Nakasora provided the first non-trivial examples of a code whose shells are $t$-designs for all weights and whose shells are $t^{\prime}$-designs for some weights with some $t^{\prime}>t$. Therefore, it is important to determine the value $t$ such that all shells of a code are $t$-designs, and no shell is a $t+1$-design, if such a $t$ exists.
@@
Using this theorem, we show that the shells of $RM_{q}(1,m)$ and $RM_{q}(1,m)^{\perp}$ are $2$-designs.
@@
Let $C=RM_{q}(1,m)$. If $q\geq 3$ and $m\geq 2$, then for any $\ell\in\mathbb{N}$, $C_{\ell}$ is not a combinatorial $3$-design.
Using Theorem 1.3, we show that $C_{\ell}$ is a $3$-$(v,k,(\lambda_{1},\lambda_{2}))$-design (see Section 2.1).
@@
By this theorem, we show that $C_{\ell}$ is a combinatorial $4$-$(v,k,(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}))$-design.
@@
This paper is organized as follows. In Section 2, we define and give some basic properties of codes, combinatorial $t$-designs, and Jacobi polynomials used in this paper. In Sections 3, 4, and 5, we show Theorems 1.1, 1.3, and 1.6, respectively.
@@ 2. Preliminaries
Let $q$ be a prime power. A $q$-ary linear code $C$ of length $n$ is a linear subspace of $\mathbb{F}_{q}^{n}$. The dual code of $C$ is the set of vectors which are orthogonal to any codewords in $C$: $C^{\perp}:=\{x\in\mathbb{F}_{q}^{n}\mid x\cdot c=0\ \mathrm{for\ all\ }c\in C\}$. For $c\in\mathbb{F}_{q}^{n}$, the weight $\mathrm{wt}(c)$ is the number of its nonzero components. The shell of a $-\ell$ is the set of codewords whose weight is $\ell$: $C_{\ell}:=\{c\in C\mid\mathrm{wt}(c)=\ell\}$.
A combinatorial $t$-design is a pair $\mathcal{D}=(\Omega,\mathcal{B})$, where $\Omega$ is a set of points of cardinality $v$, and $\mathcal{B}$ is a collection of $k$-element subsets of $\Omega$ called blocks, with the property that any $t$-element subset of $\Omega$ is contained in precisely $\lambda$ blocks. Recall [4] for the definitions of various types of designs. A pair $\mathcal{D}=(\Omega,\mathcal{B})$ is a combinatorial design with parameters $t$-$(v,k,(\lambda_{1},\dots,\lambda_{N}))$ if $\mathcal{B}$ is a collection of $k$-element subsets of $\Omega$ called blocks and every $t$-element subset of $\Omega$ is contained in $\lambda_{i}$ blocks. Note that for $N=1$, the design coincides exactly with a $t$-design.
The support of a vector $x:=(x_{1},\dots,x_{n})$, $x_{i}\in\mathbb{F}_{q}$, is the set of indices of its nonzero coordinates: $\mathrm{supp}(x)=\{i\mid x_{i}\neq 0\}$. Let $\Omega:=\{1,\dots,n\}$ and $\mathcal{B}(C_{\ell}):=\{\mathrm{supp}(x)\mid x\in C_{\ell}\}$. Then for a code $C$ of length $n$, we say that the shell $C_{\ell}$ is a combinatorial $t$-design if $(\Omega,\mathcal{B}(C_{\ell}))$ is a combinatorial $t$-design. Similarly, we say that the shell $C_{\ell}$ is a combinatorial $t$-$(v,k,(\lambda_{1},\dots,\lambda_{N}))$-design if $(\Omega,\mathcal{B}(C_{\ell}))$ is a combinatorial $t$-$(v,k,(\lambda_{1},\dots,\lambda_{N}))$-design.
@@
This fact is useful because it suffices to consider the case that $T$ contains zero.
@@ 2.3. Notation
We next introduce some notation. Let $C=RM_{q}(1,m)$, $T\subset V$, and $t=|T|$. For $c=(\lambda(x)+b)_{x\in V}\in C$ and $u\in V$, the evaluation of $c$ at $u$ is denoted by $c(u)$, which equals $\lambda(u)+b$. Let $i\in\{0,1,\dots,t\}$, $j\in\mathbb{F}_{q}$, we define
@@
$a_{i}:=|\{c\in C\setminus\mathbb{F}_{q}\mathbf{1}\mid\mathrm{wt}(c|_{T})=i\}|,$
@@
$a_{t-i}$ =|$=|\{c\in C\setminus\mathbb{F}_{q}\mathbf{1}\mid n_{0,T}(c)=i\}|$
@@
Thus, we obtain coefficients of the Jacobi polynomial by (2.1). ∎
@@
Remark.
Collorary 1.2 can be proved by $2$-transitivity of the automorphism group of $RM_{q}(1,m)$, which is the general linear homogenious group [8].
@@
Because $u_{1}$ and $u_{2}$ are not equal to 00 or each other, there exist $a,b\in\mathbb{F}_{q}\setminus\{0\}$ such that $a\neq b$ and {[$\{{}^{t}[a\ b]\}$ is a basis of $\mathrm{Im}\,A$. Hence, for any $v\in\mathrm{Im}\,A$,
@@
Then, we show that the shells of $RM_{q}(1,m)$ are not $3$-designs if $q\geq 3$ and $m\geq 2$.
@@
Let $C=RM_{q}(1,m)$. We give a proof relying on the properties of Jacobi polynomials. Let $T_{1}=\{0,u_{1},u_{2}\}\in\binom{V}{3}$, $T_{2}=\{0,v_{1},v_{2}\}\in\binom{V}{3}$, $A_{1}={}^{t}[u_{1}\ u_{2}]$, and $A_{2}={}^{t}[v_{1}\ v_{2}]$. We assume that $\mathrm{rank}\,A_{1}=2$ and $\mathrm{rank}\,A_{2}=1$. Indeed, if $q\geq 3$ and $m\geq 2$, there exist such $T_{1},T_{2}$. By Theorem 1.3,
@@
Since the coefficient of $z^{3}x^{q^{m}-\ell}y^{\ell-3}$ in $J_{C,T_{1}}-J_{C,T_{2}}$ is non-zero whenever $C_{\ell}$ is non-empty, $C_{\ell}$ is not a $3$-design. ∎
By using Theorem 2.1, we obtain
@@
Based on this equation, we conjecture the following.
Conjecture 4.3.
Let $C=RM_{q}(1,m)$. If $q\geq 3$, then for any $\ell\in\mathbb{N}$, $(C^{\perp})_{\ell}$ is not a combinatorial $3$-design.
We verified this conjecture for $q$, $m$ which satisfying $q^{2m}<10^{9}$. The computations were performed using the code available at GitHub [^1] https://github.com/yama821/GRMJacobi-paper.
Finally, we claim that the shells of $RM_{q}(1,m)$ are $3$-$(v,k,(\lambda_{1},\lambda_{2}))$-designs.
@@
(5.1) $\begin{split}b_{i,j}&=|\{c\in V\mid i\text{ elements of $Ac$ are equal to }j\}|\\ &=q^{m-\mathrm{rank}\,A}\times|\{v\in\mathrm{Im}\,A\mid i\text{ elements of $v$ are equal to }j\}|,\end{split}$
@@
(5.2) $\begin{split}b_{i,0}&=|\{c\in V\mid i-1\text{ elements of $Ac$ are equal to }0\}|\\ &=q^{m-\mathrm{rank}\,A}\times|\{v\in\mathrm{Im}\,A\mid i-1\text{ elements of $v$ are equal to }0\}|.\end{split}$
@@
By (5.1) and (5.2),
@@
Before stating the lemma under the condition of $\mathrm{rank}\,A=2$, we give a basis of $\mathrm{Im}\,A$. Since $\mathrm{Im}\,A$ is invariant under the right multiplication of an invertible matrix to $A$, we confine a basis of $\mathrm{Im}\,A$ to the following:
@@
If $\mathrm{rank}\,A=2$, we have the following.
@@
|$|\{(c_{1},c_{2})\in\mathbb{F}_{q}^{2}\mid c_{1}=j,c_{1}a+c_{2}b=j\}|$ =$=\left\{\begin{array}[]{ll}1&\text{if }b\neq 0,\\ 0&\text{if }b=0,\end{array}\right.$
|$|\{(c_{1},c_{2})\in\mathbb{F}_{q}^{2}\mid c_{2}=j,c_{1}a+c_{2}b=j\}|$ =$=\left\{\begin{array}[]{ll}1&\text{if }a\neq 0,\\ 0&\text{if }a=0,\end{array}\right.$
|$|\{(c_{1},c_{2})\in\mathbb{F}_{q}^{2}\mid c_{1}=j,c_{2}=j,c_{1}a+c_{2}b=j\}|$ =$=\left\{\begin{array}[]{ll}1&\text{if }a+b=1,\\ 0&\text{otherwise}.\end{array}\right.$
@@
Second, we compute $b_{1}$. Similarly,
$b_{1,0}=\left\{\begin{array}[]{ll}q^{m-2}(q-1)^{2}&\text{if }ab=0,\\ q^{m-2}(q-1)(q-2)&\text{otherwise}.\end{array}\right.$
@@
Finally, by using Lemma 2.4, we obtain $a_{0},\dots,a_{4}$ as follows: